# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

### v1

@@ 4 Proposed Federated Tensorized Clustering
Our proposed personalized federated learning framework integrates heat-kernel enhanced multi-view clustering with tensorized decomposition techniques at the local client level and coordinated global aggregation. The framework consists of three main components: (1) local tensorized heat-kernel enhanced multi-view fuzzy clustering, (2) tensor decomposition for efficient representation, and (3) federated aggregation with personalization mechanisms.
@@ 4.1 Tensorized Heat-Kernel Enhanced Local Clustering
Each client $L_{\ell}$ possesses multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$ where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$ represents the $h$-th view. We tensorize this data structure to capture multilinear relationships across samples, features, and views.
@@ 4.1.1 Tensor Data Organization
The multi-view data at client $\ell$ is organized into a 3-mode tensor:
@@
where $D_{[\ell]}=\max_{h}d_{h}^{(\ell)}$ is the maximum feature dimension (with zero-padding for views with fewer features). This tensorization enables:
  • Unified representation across heterogeneous views
  • Capture of cross-view feature interactions
  • Efficient tensor decomposition operations
@@ 4.1.2 Tensorized Kernel Euclidean Distance
We extend the heat-kernel enhanced distance metric to operate on tensor slices. For client $\ell$, the tensorized kernel Euclidean distance between data tensor slice $\mathcal{X}_{[\ell],i,:}^{h}$ (representing data point $i$) and cluster center tensor $\mathcal{A}_{[\ell],k,:}^{h}$ (representing cluster $k$) is:
$\text{TKED}_{\ell}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(\mathcal{X}_{[\ell],i,j}^{h}-\mathcal{A}_{[\ell],k,j}^{h}\right)^{2}\right)$ (60)
where $\delta_{[\ell]ij}^{h}$ are tensorized heat-kernel coefficients held by the $\ell$-th client, computed as:
@@ 4.1.3 Tensor Decomposition for Cluster Centers
To reduce computational complexity and enable efficient communication, we decompose the cluster center tensor using Tucker decomposition:
$\mathcal{A}_{[\ell]}\approx\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ (62)
where:
  • $\mathcal{G}_{[\ell]}\in\mathbb{R}^{c\times r_{2}\times r_{3}}$ is the core tensor
  • $\mathbf{P}_{[\ell]}\in\mathbb{R}^{c\times c}$ captures cluster relationships
  • $\mathbf{Q}_{[\ell]}\in\mathbb{R}^{D_{[\ell]}\times r_{2}}$ captures feature patterns
  • $\mathbf{R}_{[\ell]}\in\mathbb{R}^{s(\ell)\times r_{3}}$ captures view relationships
The decomposed representation reduces storage from $\mathcal{O}(c\cdot D_{[\ell]}\cdot s(\ell))$ to $\mathcal{O}(c\cdot r_{2}\cdot r_{3}+c^{2}+D_{[\ell]}\cdot r_{2}+s(\ell)\cdot r_{3})$.
4.1.4 Tensorized Local Objective Function
Each client optimizes a tensorized multi-view fuzzy clustering (T-MVFC) objective:
$J_{\ell}^{tensor}(\mathcal{U}_{[\ell]},\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]},\mathbf{v}_{[\ell]})=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h}^{\alpha}\text{ TKED}_{\ell}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (63)
@@ 4.1.4 Tensorized Local Objective Function
Theorem 2 (Tensorized Update Rules).
The alternating optimization of the tensorized objective function yields:
@@
$\mu_{jk}^{(i)}=\frac{\left(\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\text{TKED}_{i}(j,k)\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c}\left(\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\text{TKED}_{i}(j,k^{\prime})\right)^{-\frac{1}{m-1}}}$ (67)
@@
$\mathcal{G}_{i}=\arg\min_{\mathcal{G}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (68)
@@
$\mathbf{P}_{i}$ $=\arg\min_{\mathbf{P}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (69)
$\mathbf{Q}_{i}$ $=\arg\min_{\mathbf{Q}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}\times_{3}\mathbf{R}_{i})_{k,:,h}\|^{2}$ (70)
$\mathbf{R}_{i}$ $=\arg\min_{\mathbf{R}}\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\sum_{h=1}^{s_{i}}(v_{h}^{(i)})^{\alpha}\|\mathcal{X}_{i,j,:,h}-(\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R})_{k,:,h}\|^{2}$ (71)
@@
$v_{h}^{(i)}=\frac{\left(\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h)\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s_{i}}\left(\sum_{j=1}^{n_{i}}\sum_{k=1}^{c}(\mu_{jk}^{(i)})^{m}\text{TKED}_{i}(j,k,h^{\prime})\right)^{-\frac{1}{\alpha-1}}}$ (72)
Proof.
The proof follows similar principles to the non-tensorized case but operates on tensor slices. The factor matrix updates can be solved using alternating least squares (ALS) with weighted terms based on membership and view weights. ∎
4.1.5 Parallel Local Clustering Algorithm
Algorithm 1 Efficient Federated Kernel Multi-View Clustering (E-FKMVC)
Input: Local multi-view data $X_{i}$, global centers $A_{g}$, personalization parameter $\lambda_{i}$
Initialize: $A_{i}^{(0)}\leftarrow\lambda_{i}A_{g}+(1-\lambda_{i})A_{i}^{local}$, $V_{i}^{(0)}\leftarrow\frac{1}{s_{i}}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute heat-kernel coefficients
  Calculate $\delta_{j\ell}^{h,(i)}$ using Eq. 20 or 21
  // Update memberships
  Update $U_{i}^{(t)}$ using Eq. 26
  // Update cluster centers
  Update $A_{i}^{(t)}$ using Eq. 27
  // Update view weights
  Update $V_{i}^{(t)}$ using Eq. 28
  // Check convergence
  Compute $J_{i}^{local}$ using Eq. 22
until convergence or max iterations
return $U_{i}^{(t)},A_{i}^{(t)},V_{i}^{(t)}$
4.1.6 Tensorized Local Clustering Algorithm

### v2

@@ 6 Proposed Federated Tensorized Clustering
In this section, an introduction is provided to the proposed federated tensorized clustering framework, which integrates heat-kernel-enhanced multi-view clustering with tensor decomposition techniques. This integration enables the efficient management of high-dimensional multi-view data distributed across multiple clients. The central concept is to utilize tensor representations to capture the multi-modal structure of the data while leveraging heat-kernel based distance metrics to enhance clustering performance in a federated learning setting.
@@ 6.1 Tensorized Heat-Kernel Enhanced Local Clustering
Each client $\ell$ possesses multi-view data $X_{[\ell]}=\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$, where $X_{[\ell]}^{h}\in\mathbb{R}^{n(\ell)\times d_{h}^{(\ell)}}$ represents the data matrix for view $h$ with $n(\ell)$ samples and $d_{h}^{(\ell)}$ features. To effectively capture the multi-view structure, we propose a tensorized heat-kernel enhanced clustering approach. Two main components are involved: tensor representation of multi-view data and tensorized heat-kernel distance metric. Decomposition methods are then applied to represent cluster centers efficiently.
@@ 6.1.1 Tensor Data Organization
As multi-view data inherently possesses a multi-modal structure, we organize the data into a tensor format. The multi-view data from client $\ell$ is represented as a collection of matrices $\{X_{[\ell]}^{h}\}_{h=1}^{s(\ell)}$. To facilitate tensor operations, we stack these matrices along a new mode, resulting in a 3-mode tensor representation. The 3-mode tensor captures the relationships across samples, features, and views simultaneously. Mathematically, the multi-view data at client $\ell$ is organized into a 3-mode tensor as follows:
@@
where each slice $\mathcal{X}_{[\ell],:,:,h}=X_{[\ell]}^{h}$ corresponds to the data matrix of view $h$, and $D_{[\ell]}$ is the maximum feature dimension across all views for client $\ell$ ($D_{[\ell]}=\max_{h}d_{h}^{(\ell)}$). If views have varying feature dimensions, zero-padding is applied to ensure consistent tensor dimensions. Specifically, for view $h$ with $d_{h}^{(\ell)}$ features, we define:
$\mathcal{X}_{[\ell],i,j,h}=\begin{cases}X_{[\ell],i,j}^{h}&\text{if }j\leq d_{h}^{(\ell)}\\ 0&\text{if }j>d_{h}^{(\ell)}\end{cases}$ (58)
Thus, the complete tensor representation is given by:
$\mathcal{X}_{[\ell]}=\left[\mathcal{X}_{[\ell],:,:,1}|\mathcal{X}_{[\ell],:,:,2}|\ldots|\mathcal{X}_{[\ell],:,:,s(\ell)}\right]\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$ (59)
This tensorization enables a unified representation of multi-view data, allowing for efficient processing and analysis of the inherent multi-modal relationships. The tensor structure facilitates the application of tensor decomposition techniques for dimensionality reduction and feature extraction, which are crucial for effective clustering in high-dimensional spaces. Key advantages of this tensor representation include:
  • Preservation of multi-view relationships
  • Facilitation of tensor-based operations
  • Enabling advanced decomposition methods for clustering
@@ 6.1.2 Tensorized Kernel Euclidean Distance
To enable effective clustering in the tensorized multi-view space, we extend the heat-kernel enhanced distance metric to operate on tensor slices. The traditional heat-kernel enhanced distance metric (KED) is defined for vector data points. In our tensorized setting, we adapt this metric to measure distances between tensor slices corresponding to different views. The idea is to compute the distance between tensor slices while incorporating heat-kernel weighting to emphasize local structures in the data.
The tensorized kernel Euclidean distance (TKED) captures the similarity between tensor slices while incorporating the heat-kernel weighting to emphasize local structures in the data. Based on the original KED definition in Eq. 4, the TKED is formulated as follows:
$\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h},\mathcal{A}_{[\ell],k,:}^{h})=1-\exp\left(-\sum_{h=1}^{s(\ell)}\sum_{j=1}^{d_{h}^{(\ell)}}\delta_{[\ell]ij}^{h}\left(\mathcal{X}_{[\ell],i,j}^{h}-\mathcal{A}_{[\ell],k,j}^{h}\right)^{2}\right)$ (60)
where $\mathcal{A}_{[\ell]}\in\mathbb{R}^{c(\ell)\times D_{[\ell]}\times s(\ell)}$ represents the cluster center tensor for client $\ell$, with each slice $\mathcal{A}_{[\ell],k,:}^{h}$ corresponding to the cluster center for cluster $k$ in view $h$. The heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ are computed similarly to Eq. 20, but adapted for tensor slices. Specifically, the heat-kernel coefficients for the tensorized setting are defined as:
@@
As in the vector case, the heat-kernel coefficients $\delta_{[\ell]ij}^{h}$ are computed based on the normalized feature values across samples for each feature dimension $j$ in view $h$. This normalization ensures that the heat-kernel weighting effectively captures local structures in the tensorized multi-view space. The TKED metric thus provides a robust measure of similarity between tensor slices, facilitating effective clustering in the federated tensorized multi-view setting.
@@ 6.1.3 Tensor Decomposition for Cluster Centers
To reduce computational complexity and enable efficient communication, we represent cluster centers using a tensor decomposition approach. Specifically, we employ the PARAFAC2 decomposition to model the cluster center tensor $\mathcal{A}_{[\ell]}$ as:
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ are factor matrices corresponding to modes 1, 2, and 3, respectively.
PARAFAC2 is chosen for its flexibility in handling varying dimensions across one mode (samples) while maintaining consistent factor matrices in the other modes (features and views). This property is particularly advantageous in federated settings where clients may have heterogeneous data distributions. However, other tensor decomposition methods (e.g., Tucker, CP) can also be adapted based on specific application requirements.
The Tucker decomposition could be employed when the data exhibits strong multilinear interactions across all modes, while CP decomposition may be preferred for its simplicity and interpretability in scenarios with more uniform data structures. The Tucker decomposition allows for a more compact representation by capturing interactions through a core tensor, which can be beneficial when the data has complex relationships. CP decomposition, on the other hand, provides a straightforward sum of rank-one tensors, making it easier to implement and interpret in certain contexts. In this work, we also focus on Tucker decomposition as an alternative to PARAFAC2, providing flexibility in modeling diverse data characteristics across clients.
The representation cluster centers using Tucker decomposition is given by:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (63)
where $\mathcal{G}_{[\ell]}$ is the core tensor and $\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ are factor matrices corresponding to modes 1, 2, and 3, respectively.
6.2 The Proposed FedHK-PARAFAC2 Local Objective Function
The FedHK-PARAFAC2 local objective function for client $\ell$ is defined as:
$J_{FedHK-PARAFAC2}^{(\ell)}=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (64)
@@ 6.2 The Proposed FedHK-PARAFAC2 Local Objective Function
Theorem 2 (The FedHK-PARAFAC2 Update Rules).
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering at client $\ell$ are given by:
@@
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k^{\prime},:}^{h})\right)^{-\frac{1}{m-1}}}$ (68)
@@
$\mathcal{G}_{[\ell]}=\arg\min_{\mathcal{G}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (69)
View weight updates:
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h^{\prime}}-\mathcal{A}_{[\ell],k,:}^{h^{\prime}})\right)^{-\frac{1}{\alpha-1}}}$ (70)
Cluster center updates via PARAFAC2 decomposition:
@@
$\mathbf{P}_{[\ell]}$ $=\arg\min_{\mathbf{P}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (72)
$\mathbf{Q}_{[\ell]}$ $=\arg\min_{\mathbf{Q}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (73)
$\mathbf{R}_{[\ell]}$ $=\arg\min_{\mathbf{R}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{P}\times_{2}\mathbf{Q}\times_{3}\mathbf{R})_{[\ell],k,:}^{h}\|^{2}$ (74)
The Proof of Theorem 2.
By applying the method of Lagrange multipliers to the tensorized local objective in Eq. LABEL:eqn:tensor_local_obj and following similar derivation steps as in Theorem 1, we derive the update rules for membership, core tensor, factor matrices, and view weights. The key steps involve differentiating the objective with respect to each variable while enforcing the constraints, leading to the closed-form solutions presented in Eqs. 68 to 70. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1.
First, we derive the membership update by differentiating the Lagrangian with respect to $\mu_{jk}^{(i)}$ and setting it to zero, leading to Eq. 68. Next, we derive the core tensor and factor matrix updates by fixing other variables and solving the resulting least-squares problems, yielding Eqs. 69 to 74. Finally, we derive the view weight update by differentiating with respect to $v_{h}^{(i)}$ and applying the normalization constraint, resulting in Eq. 70. Finally, we verify that these updates satisfy the KKT conditions and ensure convergence under the specified conditions.
6.3 The Proposed FedHK-Tucker Local Objective Function
By implementing Tucker decomposition in Eq. 63, the FedHK-Tucker local objective function for client $\ell$ is defined as:
$J_{FedHK-Tucker}^{(\ell)}=\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})$ (75)
subject to the constraints:
$\sum_{h=1}^{s(\ell)}v_{[\ell]h}$ $=1,\quad v_{[\ell]h}\in[0,1]$ (77)
$\mathcal{A}_{[\ell]}$ $=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (78)
Theorem 3 (The FedHK-Tucker Update Rules).
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering with Tucker decomposition at client $\ell$ are given by: Membership updates:
$\mu_{[\ell]ik}^{*}=\frac{\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{m-1}}}{\sum_{k^{\prime}=1}^{c(\ell)}\left(\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k^{\prime},:}^{h})\right)^{-\frac{1}{m-1}}}$ (79)
Core tensor updates:
$\mathcal{G}_{[\ell]}=\arg\min_{\mathcal{G}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (80)
@@ 6.3 The Proposed FedHK-Tucker Local Objective Function
$v_{[\ell]h}=\frac{\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h}-\mathcal{A}_{[\ell],k,:}^{h})\right)^{-\frac{1}{\alpha-1}}}{\sum_{h^{\prime}=1}^{s(\ell)}\left(\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\text{TKED}(\mathcal{X}_{[\ell],i,:}^{h^{\prime}}-\mathcal{A}_{[\ell],k,:}^{h^{\prime}})\right)^{-\frac{1}{\alpha-1}}}$ (81)
Cluster center updates via Tucker decomposition:
$\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ (82)
Factor matrix updates:
$\mathbf{U}_{[\ell]}$ $=\arg\min_{\mathbf{U}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (83)
$\mathbf{V}_{[\ell]}$ $=\arg\min_{\mathbf{V}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (84)
$\mathbf{W}_{[\ell]}$ $=\arg\min_{\mathbf{W}}\sum_{i=1}^{n(\ell)}\sum_{k=1}^{c(\ell)}(\mu_{[\ell]ik}^{*})^{m}\sum_{h=1}^{s(\ell)}(v_{[\ell]h})^{\alpha}\|\mathcal{X}_{[\ell],i,:}^{h}-(\mathcal{G}\times_{1}\mathbf{U}\times_{2}\mathbf{V}\times_{3}\mathbf{W})_{[\ell],k,:}^{h}\|^{2}$ (85)
The Proof of Theorem 3.
The proof follows similar steps as in Theorem 2, with adjustments for the Tucker decomposition structure. By applying the method of Lagrange multipliers to the Tucker-based local objective in Eq. 75 and differentiating with respect to each variable while enforcing the constraints, we derive the closed-form solutions presented in Eqs. 79 to 81. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1. ∎
6.4 Local Clustering Algorithms
Provided below the differences between the local clustering algorithms for E-FKMVC and the tensorized FedHK-PARAFAC2/FedHK-Tucker methods. Table 1 summarizes the key distinctions.
Table 1: Comparison of Local Clustering Algorithms
Aspect E-FKMVC FedHK-PARAFAC2/FedHK-Tucker
Data Structure Multi-view matrices Multi-view tensors
Distance Metric FKED (Eq. 4) TKED (Eq. 60)
Cluster Center Representation Direct centers $A^{h}$ Decomposed tensors $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ or $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$
Update Rules Eqs. 26 - 28 Eqs. 68 - 70 or 79 - 81
Optimization Method Direct updates Alternating Least Squares (ALS) for tensor factors
As illustrated in Table 1, the E-FKMVC algorithm operates directly on multi-view data matrices using the FKED metric, while the tensorized methods utilize multi-view tensors and the TKED metric. The cluster centers in E-FKMVC are represented directly, whereas in the tensorized methods, they are represented via tensor decompositions (PARAFAC2 or Tucker). The update rules differ accordingly, with the tensorized methods employing Alternating Least Squares (ALS) for updating tensor factors.
The PARAFAC2 and Tucker decompositions in E-FKMVC enable efficient representation and updating of cluster centers in high-dimensional multi-view spaces, facilitating effective clustering in federated settings. For both tensorized methods, the local clustering algorithms involve iterative updates of memberships, tensor factors, and view weights until convergence. Table 2 summarizes the key differences between the FedHK-PARAFAC2 and FedHK-Tucker methods.
Table 2: Comparison of FedHK-PARAFAC2 and FedHK-Tucker Methods
Aspect FedHK-PARAFAC2 FedHK-Tucker
Decomposition Type PARAFAC2 Tucker
Cluster Center Representation $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$
Update Rules Eqs. 68 - 70 Eqs. 79 - 81
Optimization Method ALS for PARAFAC2 factors ALS for Tucker factors
As can be seen in Table 2, the primary difference between the two methods lies in the type of tensor decomposition employed (PARAFAC2 vs. Tucker) and the corresponding representation of cluster centers. The update rules and optimization methods are adapted accordingly to suit the specific decomposition framework used in each method. See Algorithm 2 and Algorithm 3 for the detailed local clustering procedures for FedHK-PARAFAC2 and FedHK-Tucker, respectively.
ALS Optimization for PARAFAC2 Factors
The Alternating Least Squares (ALS) optimization for updating the PARAFAC2 factors involves iteratively optimizing each factor matrix while keeping the others fixed. The update steps for the factor matrices $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ are derived by solving least-squares problems based on the current estimates of the other factors and the core tensor $\mathcal{G}_{[\ell]}$. This process continues until convergence, ensuring that the cluster center tensor $\mathcal{A}_{[\ell]}$ is accurately represented in the decomposed form.
ALS Optimization for Tucker Factors
Similar to the PARAFAC2 case, the ALS optimization for Tucker factors involves iteratively updating each factor matrix $\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ while keeping the others fixed. The updates are obtained by solving least-squares problems that minimize the reconstruction error of the cluster center tensor $\mathcal{A}_{[\ell]}$ based on the current estimates of the core tensor $\mathcal{G}_{[\ell]}$ and the other factor matrices. This iterative process continues until convergence, ensuring an accurate representation of the cluster centers in the Tucker decomposed form.
For both tensorized methods, the ALS optimization provides an efficient means of updating the tensor factors, enabling effective clustering in high-dimensional multi-view spaces while leveraging the benefits of tensor decomposition. Please refer to previous sections for the specific update equations and procedures involved in the ALS optimization for both PARAFAC2 and Tucker decompositions.